OpenStudy (anonymous):

use the quadratic formula to solve, -2x^2+9x+3=0

OpenStudy (helder_edwin):

given the equation \(ax^2+bx+c=0\) the quadratic formula says \[\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (helder_edwin):

in your example a=?, b=?, and c=?